OpenStudy (anonymous):

a building in a city has a rectangler base. the length of the base measures 60 ft less than three times the width. the perimeter of this base is 880. what are the dimensions of the base? @DebbieG

OpenStudy (debbieg):

You had almost the identical problem here: http://openstudy.com/users/anton99#/updates/52310eb7e4b03eb771a2610b and it was pretty much done for you. Did you try to work through that one and understand it? And have you tried to sketch this one?

OpenStudy (anonymous):

yeah i tried except its 3 times

OpenStudy (anonymous):

i keep getting the wrong answer

OpenStudy (debbieg):



OpenStudy (debbieg):

" the length of the base measures 60 ft less than three times the width." How can you put that into a mathematical expression? use W for width. Then L=??? (in terms of W)

OpenStudy (anonymous):

yeah how can u

OpenStudy (debbieg):

If W is width, then what is "3 times the width"?

OpenStudy (anonymous):

3w

OpenStudy (debbieg):

Good. Now, what is: "60 ft less than three times the width."

OpenStudy (anonymous):

3w-60

OpenStudy (debbieg):

See? You understand more than you think you do, you just need to think it through one step at a time! :) Now you have: L=3W - 60 That "models" the information in the problem: " the length of the base measures 60 ft less than three times the width." Now we will do something with that, by using the perimeter information:

OpenStudy (debbieg):

You are told that perimeter is 880. and there is a STANDARD formula for perimeter of a rectangle (you should know this!) P=2L + 2W Now, you have L=3W - 60, and what THAT does is allow us to take the perimeter equation (which has TWO unknowns, L and W), and make a substitution so that we end up with an equation with just ONE variable - and that, we can solve.

OpenStudy (debbieg):

So we have to SUBSTITUTE the expression for length, L=3W - 60, into the perimeter equation - just "drop it in" where we have the L! And also, put in the "known value" of perimeter, for P: P=2L + 2W 880=2(3W - 60) + 2W

OpenStudy (debbieg):

Are you with me? Stop me if you don't understand something to this point.

OpenStudy (debbieg):

ask questions.

OpenStudy (anonymous):

125 = w

OpenStudy (debbieg):

Exellent!! That's correct! And now, all you need is to get length... guess what? you have a conveeeenient formula for length, in terms of W: L=3W - 60 So just take that W that you just figured out, and plug & chug in the length formula, and tah-dah, you have length and width. :)

OpenStudy (anonymous):

275?

OpenStudy (debbieg):

For length? Not quite.....

OpenStudy (debbieg):

L=3W - 60 plug in W=125....

OpenStudy (anonymous):

GOT IT